k B Busey Bank - Oniline DWaves Bloomington ostpay 8 Em of 24 -4 MR uestion 4 of 27 1 Points Two cities each had a population of 1,000,000. The death rate from all causes in Desert City was 500 (or 50 per 100,000). The death rate from all causes in Sun City was 800 (or 80 per 100,000). Cancer deaths in both cities were 40 deaths (or 4 per 100,000) per city. The risk of cancer-caused death for both cities was the same. Calculate the PMR (proportional mortality ratio) for cancer in each city. OA. Desert City 0.0008; Sun City 0.0005 B. Desert City 0.004%; Sun City 0.004% 0 c. Desert City 0.08%; Sun City 0.05% D. Desert City 8%; Sun City 5% eset Selection Mark for Review Whats This?
Answer #1

Proportional mortality ratio=

Mortality due to specific cause during a time period/mortality due to all cause x 100

For dessert city = 40/500 x 100 = 8%

For suncity = 40/800 x 100 =5%

So,Option D is the correct answer.